What is the OH+ of a solution whose POH =4.70

1
Expert's answer
2022-05-24T06:06:03-0400

Solution:

We can convert between pOH and [OH] using the following equation:

pOH = −log[OH]

[OH] = 10−pOH = 10−4.70 = 2.00×10−5

[OH] = 2.00×10−5 M


For any aqueous solution at 25C:

pH + pOH = 14

pH = 14 − pOH = 14 − 4.7 = 9.3

pH = 9.3


We can convert between pH and [H+] using the following equation:

pH = −log[H+]

[H+] = 10−pH = 10−9.3 = 5.01×10−10

[H+] = 5.01×10−10 M


Answer:

The hydroxide ion (OH) concentration is 2.00×10−5 M

The hydrogen ion (H+) concentration is 5.01×10−10 M